When 25.0 mL of 0.700 mol/L NaOH was mixed in a calorimeter with 25.0 mL of 0.700 mol/L HCl, both initially at 20.0 °C, the temperature increased to 22.1 °C. The heat capacity of the calorimeter is 279 J/°C. What is the enthalpy of neutralization in kJ/ mole of HCl?

Since the solutions are mostly water, the solutions are assumed to have a density of 1.0 g/mL and a specific heat of 4.18 J/g°C. 

Select one:

a. −1020

b. -58.6 

c. −5856

d. 58.6
Answer #1

no of moles of HCl   = molarity * volume in L

                                = 0.7*0.025   = 0.0175 moles

total volume of solution = 25 +25 = 50ml

mass of solution = volume * desnity

                             = 50*1 = 50g

The heat involves are

heat from neutralization +heat to warm solution + heat to warm calorimeter = 0

                                    q1 + q2 + q3 = 0

                                     n\DeltaH = mC\DeltaT + C\DeltaT   = 0

                                      0.0175mole*\DeltaH = 50*4.18*(22.1-20) + 279*(22.1-20) = 0

                   \Delta H   = -58560J/mole

                             = -58.56KJ/mole

round off -58.6KJ/mole

b. -58.6   >>>>answer